# Cosmological dynamics of interacting dark energy and dark matter in $f(Q)$ gravity

### v1

@@ III Autonomous dynamical system of interacting dark energy and dark matter
$x=\frac{f}{12H^{2}\,F},\,\,\,\,\,y=\frac{\rho_{r}}{6H^{2}\,F}.$ (9)
Invoking the above dimensionless variables, it is straightforward to derive the autonomous equations which play a key role for studying the dynamical system for the interacting dark energy and dark matter.
@@ III.2 Case II : $\mathcal{U}=\frac{\alpha}{3H}\rho_{m}\,\rho_{de}$
Finally, we define the equation of state and deceleration parameters corresponding to the dimensionless variables to check the appropriate acceleration expansion of the universe, that is
@@
Both $\omega$ and $q$ represent different aspects of cosmic evolution, each uniquely influencing the large-scale structure of the universe. Grasping the significance of these parameters is essential for studying the effects of dark energy across various stages of cosmic evolution. Moreover, these parameters aid in comparing and differentiating between various dark energy models, each employing distinct mechanisms to drive cosmic acceleration.
@@ IV Model:- $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$
In the previous section, we derived autonomous dynamical systems, and our primary objective is to investigate and analyze them. To accomplish this goal, we will focus on the designated power-law $f(Q)$ model, which has the form $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$, where $\gamma$ and $n$ are free model parameters.
@@
Now, we will find the critical points and corresponding eigenvalues for both the dynamical systems. Studying the critical points and assessing their stability is crucial for a comprehensive understanding of critical aspects in cosmic evolution driven by dark energy and dark matter interactions, as explored within the scope of this study. The properties of the dynamical system depend on the values of the constants $\alpha$ and $n$.
@@ IV.1 For case I
The fixed points $(x,y)$ of the general dynamical system are the following:
$P_{1}:(1,0),\,\,\,\,P_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right),$ (19)
$P_{3}:\left(\frac{1}{2n+\alpha(2n-1)},0\right).$ (20)
Furthermore, the corresponding eigenvalues for each fixed point are given by
$P_{1}:\{-4,-3(\alpha+1)\},\,\,\,\,P_{2}:\{4,1-3\alpha\},$ (21)
$P_{3}:\{3(\alpha+1),3\alpha-1\}.$ (22)
Let us now focus on the dynamics of each critical point and its features in the further subsections.
@@ IV.1.1 $P_{1}$: the de Sitter fixed point
This critical point is independent of both the model and coupling parameters. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are
@@
At this point, both dark matter and radiation are absent, and the universe is dominated by dark energy. The equation of state (EoS) at this fixed point indicates that the universe is experiencing an accelerated expansion. To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point. The eigenvalues for the fixed point $P_{1}$ read
$\{-4,\,\,\,-3(\alpha+1)\}.$ (24)
The stability behavior of this fixed point (which depends on $\alpha$) is obtained as:
@@ IV.1.2 $P_{2}$: Non-metricity dominated fixed point
The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters are obtained as
@@
The conditions for an accelerating universe are $n<0$ or $n>2$. When $n<0$, the universe exhibits phantom-like behavior, i.e., $\omega<-1$. When $n>2$, the universe exhibits quintessence-like behavior, i.e., $-1<\omega<-1/3$. In this fixed-point solution, a matter-dominated era occurs for $n=4/3$ and a radiation-dominated era for $n=1$. The eigenvalues for the fixed point $P_{2}$ are as follows:
{4,   $\{4,\,\,\,1-3\alpha\}.$ (26)
The stability behavior for this particular fixed point can be obtained for different $\alpha$, given by
@@ IV.1.3 $P_{3}$: scaling solution fixed point
It is worth noting that our result modifies the scaling solution of $f(Q)$ gravity with the interacting dark energy parameter $\alpha$. The fixed point $P_{3}$ reads
@@
The properties of this critical point are influenced by both the model parameter $n$ and the coupling parameter $\alpha$. Corresponding to this fixed point, the matter density, dark energy density, and effective EoS parameters read as
$\Omega_{m}=\frac{1}{\alpha-2(\alpha+1)n}+1,\,\,\,\,\Omega_{DE}=\frac{1}{2n+\alpha(2n-1)},$
and
$\omega=\frac{3\alpha-2(\alpha+1)n+2}{2(\alpha+1)n-\alpha}.$
@@
Figure 1: In this figure, we show the relationship between the model parameter $n$ and the interacting parameter $\alpha$ for the accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
In this fixed point solution, we obtained a matter-dominated era for $n=\frac{3\alpha+2}{2(\alpha+1)}$, $1+\alpha\neq 0$, and radiation-dominated era for $n=\frac{5\alpha+3}{4(\alpha+1)}$, $1+\alpha\neq 0$.
The eigenvalues for the fixed point $P_{3}$ read as
@@
We determine the stability of this fixed point for various ranges of $\alpha$. The ranges are as follows:
  • Stable node for $\alpha<-1$
  • Saddle for $-1<\alpha<1/3$
@@
In Figs. 2 and 3, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. This is done using the numerical solution of the dynamical system with the initial conditions $x(0)=0.7$ and $y(0)=0.00005$. The motivation behind the choice of initial conditions is as follows: $x$ represents the dark energy density, which currently has a value of $0.7$. Similarly, $y$ represents the radiation density, which currently has a value of $0.00005$.
@@
Figure 2: The plot on the left displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the positive coupling parameter, which can transfer energy from dark matter to dark energy.
@@
Figure 3: The left panel displays density parameters for matter ($\Omega_{m}$), dark energy ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the equation of state parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from dark energy to dark matter.
@@
For $n=3/2$ and $\alpha=1/2$ in Fig. 2, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from dark matter to dark energy. This figure shows that the universe was dominated by matter in the early stages and will be dominated by dark energy in later stages. Currently, the universe is dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$. For these values, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle-node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration.
@@
In Fig. 3, with $n=3/2$ and $\alpha=-2$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, that means energy transfers from dark energy to dark matter. This figure shows that the universe was dominated by dark energy in the early stages and will be dominated by dark matter at later stages. Currently, the universe remains dominated by dark energy, with parameter values $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, and $\Omega_{DE}=0.7$, $q_{0}=-0.563$, and $\omega_{0}=-0.71$. Here, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle-node, and $P_{2}$ is an unstable node. Fig. 4 shows the phase space trajectories of the fixed points $P_{1}$, $P_{2}$, and $P_{3}$.
Figure 4: Phase plot of $x$ vs $y$. The upper phase plot shows the stable behavior of $P_{1}$ for $\alpha>0$, and the lower phase plot shows the stable behavior of $P_{3}$ for $\alpha<-1$.
@@ IV.2 For case II
Next, the second case in $f(Q)$ gravity is also investigated in detail for the interacting dark energy system. The models have been systematically constructed with possible ranges of their model parameters derived from the dynamical system analysis within the interacting dark energy framework. It is important to examine how the interaction affects the cosmological viability of the $f(Q)$ gravity models.
The fixed points $(x,y)$ of the dynamical system are the following:
$R_{1}:(1,0),\,\,\,\,R_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right),R_{3}:\left(\frac{1}{\alpha},\frac{-3\alpha+8n-1}{\alpha}\right),$
$R_{4}:\left(\frac{-\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right),$ (28)
$R_{5}:\left(\frac{\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right).$
The corresponding eigenvalues are
$R_{1}:\{-4,-3-\alpha\},\,\,\,\,R_{2}:\{4,1-\frac{\alpha}{2n}\},$
$R_{3}:\{\mathcal{A}_{1}-\frac{4n}{\alpha}+4,-\mathcal{A}_{1}-\frac{4n}{\alpha}+4\},$
$R_{4}:\left\{\frac{5n-1+\mathcal{A}_{2}}{1-2n},\frac{-2\mathcal{A}_{2}^{2}-2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha}\right\},$
$R_{5}:\left\{\frac{-2\mathcal{A}_{2}^{2}+2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha},\frac{1-5n+\mathcal{A}_{2}}{2n-1}\right\}.$
where $\mathcal{A}_{1}=\frac{2\sqrt{(2n-1)\left(3(\alpha-1)\alpha^{2}+8n^{3}-4n^{2}-6(\alpha-1)\alpha n\right)}}{\alpha-2\alpha n}$ and $\mathcal{A}_{2}=\sqrt{9n^{2}-3\alpha+6n\alpha}$.
@@ IV.2.1 $R_{1}$: the de Sitter fixed point
Corresponding to fixed point $R_{1}:(x,y)=(1,0)$, the matter density, dark energy density, and effective EoS parameters are
@@
At this point, both dark matter and radiation are absent, and the universe is dominated by dark energy. The EoS at this fixed point indicates that the universe is experiencing an accelerated expansion. To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point. The eigenvalues for the fixed point $R_{1}$ read
$\{{-4,\,\,\,-\alpha-3\}}.$ (30)

### v2

@@ III Autonomous dynamical system of interacting DE and DM
$x=\frac{f}{12H^{2}\,f_{Q}},\,\,\,\,\,y=\frac{\rho_{r}}{6H^{2}\,f_{Q}}.$ (13)
Invoking the above dimensionless variables, it is straightforward to derive the autonomous equations which play a key role for studying the dynamical system for the interacting DE and DM.
@@ III.2 Case II : $\mathcal{U}=\frac{\alpha}{3H}\rho_{m}\,\rho_{DE}$
Finally, we define the EoS and deceleration parameters corresponding to the dimensionless variables to check the appropriate acceleration expansion of the universe, that is
@@
Both $\omega$ and $q$ represent different aspects of cosmic evolution, each uniquely influencing the large-scale structure of the universe. Grasping the significance of these parameters is essential for studying the effects of DE across various stages of cosmic evolution. Moreover, these parameters aid in comparing and differentiating between various DE models, each employing distinct mechanisms to drive cosmic acceleration.
@@ IV Model:- $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$
In the previous section, we derived autonomous dynamical systems, and our primary objective is to investigate and analyze them. To accomplish this goal, we will focus on the designated power-law $f(Q)$ model, which has the form $f(Q)=6\gamma\,H_{0}^{2}\left(\frac{Q}{Q_{0}}\right)^{n}$, where $\gamma$, $n$ are free model parameters and $Q_{0}=6H_{0}^{2}$.
@@
We will now derive the critical points of the system by setting $dx/dN=dy/dN=0$ and determine the corresponding eigenvalues for both dynamical systems. The critical points, which represent the solutions of the dynamical system, provide an initial qualitative understanding of the phase space. As discussed below, these points can be classified based on their stability properties. In the absence of singularities or strange attractors, the trajectories of $x(N)$ and $y(N)$, which are generally obtained numerically, tend to evolve from unstable fixed points to stable fixed points, possibly passing through intermediate saddle points.
Analyzing the critical points and evaluating their stability is essential for a thorough understanding of the critical aspects of cosmic evolution driven by interactions between DE and DM, as explored in this study. Furthermore, the properties of the dynamical system depend significantly on the values of the constants $\alpha$ and $n$.
@@ IV.1 For case I
The fixed points $(x,y)$ of the general dynamical system are described in table 1. Let us now focus on the dynamics of each critical point and its features in the following subsections.
Critical Points $\Omega_{m}$ $\Omega_{DE}$ $w$ Eigenvalues Stability
$P_{1}:(1,0)$ 00 $1$ $-1$ $\{-4,\,\,\,-3(\alpha+1)\}$
Stable node for $\alpha>-1$,
Saddle for $\alpha<-1$,
Non-hyperbolic for $\alpha=-1$.
$P_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right)$ 00 $\frac{1}{2n}$ $-1+\frac{4}{3n}$$\{4,1-3\alpha\}$
Saddle node for $\alpha>1/3$,
Unstable for $\alpha<1/3$,
Non-hyperbolic for $\alpha=1/3$.
$P_{3}:\left(\frac{1}{2n+\alpha(2n-1)},0\right)$ $\frac{1}{\alpha-2(\alpha+1)n}+1$ $\frac{1}{2n+\alpha(2n-1)}$ $\frac{3\alpha-2(\alpha+1)n+2}{2(\alpha+1)n-\alpha}$$\{3(\alpha+1),3\alpha-1\}$
Stable node for $\alpha<-1$,
Saddle for $-1<\alpha<1/3$,
Unstable node for $\alpha>1/3$.
Table 1: This table summarizes the analysis of the critical points for Case I.
@@ IV.1.1 $P_{1}$: the de Sitter fixed point
This critical point is independent of both the model and coupling parameters. Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are
@@
At this point, both DM and radiation are absent, and the universe is dominated by DE. The EoS at this fixed point indicates that the universe is experiencing an accelerated expansion. To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point. The eigenvalues for the fixed point $P_{1}$ read $\{-4,\,\,\,-3(\alpha+1)\}$. The stability behavior of this fixed point (which depends on $\alpha$) is obtained as
@@ IV.1.2 $P_{2}$: Non-metricity dominated fixed point
The characteristics of this critical point solely depend on the model parameter $n$. Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters are obtained as
@@
The conditions for an accelerating universe are $n<0$ or $n>2$. When $n<0$, the universe exhibits phantom-like behavior, i.e., $\omega<-1$. When $n>2$, the universe exhibits quintessence-like behavior, i.e., $-1<\omega<-1/3$. In this fixed-point solution, a matter-dominated era occurs for $n=4/3$ and a radiation-dominated era for $n=1$. The eigenvalues for the fixed point $P_{2}$ are as follows {4,   $\{4,\,\,\,1-3\alpha\}$. The stability behavior for this particular fixed point can be obtained for different $\alpha$, given by
@@ IV.1.3 $P_{3}$: scaling solution fixed point
It is worth noting that our result modifies the scaling solution of $f(Q)$ gravity with the interacting DE parameter $\alpha$. The fixed point $P_{3}$ reads
@@
The properties of this critical point are influenced by both the model parameter $n$ and the coupling parameter $\alpha$. Corresponding to this fixed point, the matter density, DE density, and effective EoS parameters read as
$\Omega_{m}=\frac{1}{\alpha-2(\alpha+1)n}+1,\quad\Omega_{DE}=\frac{1}{2n+\alpha(2n-1)},\quad\omega=\frac{3\alpha-2(\alpha+1)n+2}{2(\alpha+1)n-\alpha}.$
@@
Figure 1: The plot illustrates the relationship between the model parameter $n$ and the interaction parameter $\alpha$ for an accelerating universe. The light blue region represents the quintessence-like behavior of an accelerating universe (i.e., $-1<\omega<-1/3$), while the golden region represents the phantom-like behavior of an accelerating universe (i.e., $\omega<-1$).
@@
  • Stable node for $\alpha<-1$,
  • Saddle for $-1<\alpha<1/3$,
@@
In Figs. 2 and 3, we illustrate the evolution of $\Omega_{m}$, $\Omega_{r}$, $\Omega_{DE}$, $q$, and $\omega$ for different values of the model parameter $n$ and the interaction parameter $\alpha$. This is done using the numerical solution of the dynamical system with the initial conditions $x(0)=0.7$ and $y(0)=0.00005$. The motivation behind the choice of initial conditions is as follows: $x$ represents the DE density, which currently has a value of $0.7$. Similarly, $y$ represents the radiation density, which currently has a value of $0.00005$.
@@
Figure 2: The left panel shows the density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift ($z$). The right panel displays the deceleration parameter ($q$) and the equation of state parameter ($\omega$) as functions of redshift ($z$). Together, these two panels illustrate the trajectory for a positive coupling parameter, which can facilitate the transfer of energy from DM to DE.
@@
Figure 3: The left panel displays density parameters for matter ($\Omega_{m}$), DE ($\Omega_{DE}$), and radiation ($\Omega_{r}$) as functions of redshift $z$. The right panel shows the deceleration parameter $q$ and the EoS parameter $\omega$ as functions of redshift $z$. These two panels illustrate the trajectory for the negative coupling parameter, which can transfer energy from DE to DM.
@@
For $n=3/2$ and $\alpha=1/2$ in Fig. 2, $\alpha>0$ indicates that the coupling term $\mathcal{U}>0$, signifying energy transfer from DM to DE. This figure shows that the universe was dominated by matter in the early stages and will be dominated by DE in later stages. Currently, the universe is dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.55$, and $\omega_{0}=-0.70$. For these values, fixed point $P_{1}$ is stable and represents the de Sitter acceleration solution, while fixed point $P_{2}$ is a saddle-node, and $P_{3}$ is an unstable node that cannot demonstrate universal acceleration.
@@
In Fig. 3, with $n=3/2$ and $\alpha=-2$, $\alpha<0$ indicates that the coupling term $\mathcal{U}<0$, that means energy transfers from DE to DM. This figure shows that the universe was dominated by DE in the early stages and will be dominated by DM at later stages. Currently, the universe remains dominated by DE, with parameters value $\Omega_{m}=0.3$, $\Omega_{r}=0.00005$, $\Omega_{DE}=0.7$, $q_{0}=-0.563$, and $\omega_{0}=-0.71$. Here, fixed point $P_{3}$ is stable and exhibits acceleration for the early and present universe but fails to show acceleration for late times, while fixed point $P_{1}$ is a saddle-node, and $P_{2}$ is an unstable node. Fig. 4 shows the phase space trajectories of the fixed points $P_{1}$, $P_{2}$, and $P_{3}$.
Figure 4: Phase space of $x$ vs $y$. The left phase plot illustrates the stable behavior of $P_{1}$ when $\alpha>0$, while the right phase plot showcases the stable behavior of $P_{3}$ for $\alpha<-1$.
@@ IV.2 For case II
Next, the second case in $f(Q)$ gravity is also investigated in detail for the interacting DE system. The models have been systematically constructed with possible ranges of their model parameters derived from the dynamical system analysis within the interacting DE framework. It is important to examine how the interaction affects the cosmological viability of the $f(Q)$ gravity models.
The fixed points $(x,y)$ of the dynamical system are described in table 2.
Critical Points $\Omega_{m}$ $\Omega_{DE}$ $w$ Eigenvalues Stability
$R_{1}:(1,0)$ 00 $1$ $-1$ $\{-4,-3-\alpha\}$
Stable node for $\alpha>-3$,
Saddle for $\alpha<-3$,
Non-hyperbolic for $\alpha=-3$.
$R_{2}:\left(\frac{1}{2n},\frac{2n-1}{2n}\right)$ 00 $\frac{1}{2n}$ $-1+\frac{4}{3n}$$\{4,1-\frac{\alpha}{2n}\}$
Saddle node for $\alpha>2n$,
Unstable for $\alpha<2n$,
Non-hyperbolic for $\alpha=2n$.
$R_{3}:\left(\frac{1}{\alpha},\frac{-3\alpha+8n-1}{\alpha}\right)$ $4-\frac{8n}{\alpha}$ $\frac{1}{\alpha}$ $\frac{8}{3\alpha}-1$$\{\mathcal{A}_{1}-\frac{4n}{\alpha}+4,-\mathcal{A}_{1}-\frac{4n}{\alpha}+4\}$
The stability conditions
are shown in Fig. 5
$R_{4}:\left(\frac{-\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right)$ Sec.IV.2.4 IV.2.4 IV.2.4 $\left\{\frac{5n-1+\mathcal{A}_{2}}{1-2n},\frac{-2\mathcal{A}_{2}^{2}-2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha}\right\}$
The stability conditions
are shown in Fig. 5
$R_{5}:\left(\frac{\sqrt{-3\alpha+9n^{2}+6\alpha n}-3n}{2\alpha n-\alpha},0\right)$ Sec.IV.2.5 IV.2.5 IV.2.5 $\left\{\frac{-2\mathcal{A}_{2}^{2}+2(3n-\alpha+2n\alpha)\mathcal{A}_{2}}{(2n-1)^{2}\alpha},\frac{1-5n+\mathcal{A}_{2}}{2n-1}\right\}$
The stability conditions
are shown in Fig. 5
Table 2: This figure summarizes the results of the analysis of the critical points for Case II. The critical points $R_{4}$ and $R_{5}$, however, are difficult to present comprehensively in a table format. Detailed analyses of these points can be found directly in Sections IV.2.4 and IV.2.5.
The expressions of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ in the table are given by $\mathcal{A}_{1}=\frac{2\sqrt{(2n-1)\left(3(\alpha-1)\alpha^{2}+8n^{3}-4n^{2}-6(\alpha-1)\alpha n\right)}}{\alpha-2\alpha n}$ and $\mathcal{A}_{2}=\sqrt{9n^{2}-3\alpha+6n\alpha}$.
@@ IV.2.1 $R_{1}$: the de Sitter fixed point
Corresponding to fixed point $R_{1}:(x,y)=(1,0)$, the matter density, DE density, and effective EoS parameters are
@@
At this point, both DM and radiation are absent, and the universe is dominated by DE. The EoS at this fixed point indicates that the universe is experiencing an accelerated expansion. To study the stability of this fixed point, we can analyze the eigenvalues of the Jacobian matrix, which describe the behavior of the fixed point. The eigenvalues for the fixed point $R_{1}$ are $\{{-4,\,\,\,-\alpha-3\}}$.